# Virne: A Comprehensive Benchmark for Deep RL-based Network Resource Allocation in NFV

### v1

@@ A.2 Extensions in Emerging Networks
In modern data center networks, computing resources are often heterogeneous, meaning physical nodes may have varying capacities in terms of CPU, GPU, memory, etc. To account for this, we extend the basic model by incorporating heterogeneous resources, where both virtual and physic nodes are associated with a set of computing resources $\mathcal{C}$. Thus, a physical node $n_{p}$ for placing virtual node $n_{v}$ must have enough resources across all types, i.e., $C(n_{v})\leq C(n_{p}),\forall n_{v}\in\mathcal{N}_{v},C\in\mathcal{C},n_{p}=f_{\mathcal{N}}(n_{v})$.
@@ Appendix B Related Work
As network scenarios become increasingly diverse and complex (e.g., cloud, edge, 5G), many studies explore NFV-RA in emerging networks while addressing additional factors (e.g., latency, energy, heterogeneity). Despite these variations, many of these studies share a common core methodology. In this section, we focus on the methodological aspects and review the development of key approaches in NFV-RA. Then, we describe existing benchmarks to emphasize the gap between Virne. Finally, we compare NFV-RA with other COPs, highlighting the need for specific designs.
@@
[26] Edge / A2C LSTM, Attention Not available
[38] Edge Energy, Latency PG CNN Not available
[39] 5G Slicing Security DQN Heterogeneous GCN Not available
[40] 5G Slicing Latency PPO MLP Not available
[41] Internet of Things Latency PPO GNN Not available
[42] Internet of Things Latency DQN CNN Not available
[43] Satellite Latency DQN MLP Not available
[27] Data Plane Latency A3C GCN Not available
[44] Space-Air-Ground Latency PG CNN Not available
[10] Cloud / MCTS / VNE-Sim
[12] Cloud / PG RNN Not available
[45] Cloud / PG GCN Not available
[46] Cloud Security PG CNN Not available
[47] Cloud Energy PG LSTM; Attention Self-implemented
[15] Cloud / A3C GCN, GRU VNE-Sim
[48] Cloud Energy A2C Attention Virne
[49] Cloud / PG MLP Not available
[16] Cloud / PG CNN Not available
[14] Cloud / A3C GCN VNE-Sim
[11] Cloud Latency PG MLP Self-implemented
[13] Cloud / A3C GCN, GRU Virne
[18] Cloud / PPO GNN Virne
[19] Cloud Heterogeneity PPO Cross-GCN Virne
[17] Cloud / PPO Edge-aware GAT Virne
[50] Cloud, Edge Latency PPO Heterogeneous GAT Virne
@@ B.1 Traditional NFV-RA Algorithms
Earlier studies for the NFV-RA problem employed exact solvers, such as integer linear programming (ILP)[6] and mixed integer programming (MIP)[5], to find optimal solutions. However, their high computational complexity makes them impractical for real-world scenarios, especially in larger networks where dynamic service requests require rapid solutions. To address these limitations, heuristic-based methods emerged as alternatives, offering suboptimal yet computationally efficient solutions. Among these, node-ranking strategies stand out as a prominent approach [7, 8, 51, 9]. These strategies construct solutions by prioritizing the virtual and physical nodes based on specific metrics to guide the node mapping process, followed by link mapping. For instance, Node Resource Management (NRM) [8] ranks nodes by weighting both node and link resources, while the Node Essentiality Assessment (NEA) [9] incorporates topology connectivity into decision-making. Although reducing computational overhead, they rely heavily on manual design with domain-specific knowledge and lack adaptation to diverse scenarios, which can lead to inferior results in complex NFV-RA requirements. Furthermore, meta-heuristics have been adopted to solve NFV-RA by modeling the solution search as an evolutionary process [52, 53, 54]. These algorithms, such as Genetic Algorithms (GA) [55, 52] and Particle Swarm Optimization (PSO) [56, 57], explore the solution space by iteratively generating and evolving a population of candidate solutions. However, they are computationally expensive and hyperparameter-sensitive, limiting their practicality.
@@ B.2 RL-based NFV-RA Algorithms
In recent years, machine learning (ML)-based methods have gained prominence in solving NFV-RA problems due to their superior performance and adaptability to dynamic network conditions [35]. Several works [58, 59] have explored predictive models trained on high-quality datasets to forecast the future service requests. However, acquiring labeled, high-quality datasets for large-scale and unseen network scenarios remains impractical, limiting their applicability. More dominantly, RL has demonstrated its promise for NFV-RA tasks [10, 11, 12, 13, 14, 15, 16, 17, 18, 19], mainly due to its label-free nature and adaptation to handle dynamics of network. RL-based methods model NFV-RA as a Markov Decision Process (MDP), allowing an agent to learn optimal policies through iterative interactions with the environment. In general, we unified existing RL-based methods for NFV-RA under a framework with three core components: MDP modeling, policy architecture, and training methods. These methods conceptualize the node mapping allocation process as a sequential decision-making task, where physical nodes are incrementally selected to host VNFs. To execute these decisions, various policy network architectures are designed to represent the network state and generate corresponding actions. Then, they leverage a specific RL method to optimize the policy network with collected experience during interactions. For example, [11] utilized a multilayer perceptron (MLP) and trained it using the policy gradient (PG) algorithm, while the work [14] combined MLPs with graph convolutional networks (GCNs) and employed the asynchronous advantage actor-critic (A3C) algorithm. While RL-based methods offer notable advantages in terms of performance and adaptability, there is a lack of in-depth analysis of the impact of specific model design choices and consistent performance comparison. Moreover, systematically identifying the remaining significant challenges in RL-based NFV-RA is critical to open opportunities for future exploration and innovation.
@@ B.3 Existing NFV-RA Benchmarks
For the NFV-RA problem, several benchmarks have been developed for simulation and evaluation. As summarized in Table 1, the most notable existing benchmarks include: VNE-Sim [^2] https://tehreemf.wixsite.com/vne-sim [21], ALEVIN  [^3] https://sourceforge.net/p/alevin/wiki/home/ [22], ALib [^4] https://github.com/vnep-approx/alib [22], SFCSim [^5] https://github.com/SFCSim/SFCSim [24], and Iflye [^6] https://github.com/Echtzeitsysteme/iflye [25]. For instance, VNESim[21] supports three heuristic algorithms and focuses exclusively on cloud-based simulations, evaluating solutions based on their effectiveness. Similarly, SFCSim [24], incorporates three additional heuristic methods, continuing the trend of focusing on cloud-based scenarios However, most of these benchmarks are limited to cloud-based simulations and lack the flexibility to accommodate more modern network scenarios, such as edge computing or 5G environments. Moreover, these benchmarks generally implement only a few exact and heuristic algorithms and focus mainly on effectiveness evaluation. These limitations highlight the need for more comprehensive benchmarks that can assess algorithms across a broader range of network environments and evaluation perspectives.
@@ B.4 RL for Combinatorial Optimization
RL has gained significant attention for solving Combinatorial Optimization Problems (COPs), where the goal is to learn high-quality solutions [60]. These problems span various domains, including the Traveling Salesman Problem (TSP) [61], Vehicle Routing Problem (VRP) [62], and bin packing [63]. RL-based approaches to solving these problems generally fall into two categories: construction methods that build a solution incrementally from scratch and improvement methods that start with an initial solution and iteratively refine it. Compared with these classic COPs, the NFV-RA problem presents unique challenges, highlighting the need for specialization of RL approach design. First, NFV-RA requires real-time decision-making to meet strict network service requirements. NFV-RA algorithms must execute decisions almost immediately, making construction methods more suitable to incrementally build solutions within time constraints. Second, NFV-RA operates in highly dynamic environments where both service demands and network resources fluctuate in real time. This variability requires RL methods that can adapt to changes in network topologies, resource availability, and demand patterns. Lastly, the state space of NFV-RA is highly complex, involving intricate interactions such as cross-graph mapping and bandwidth-constrained routing. These factors require RL methods to navigate a multidimensional decision space, accounting for diverse resource types and connectivity requirements.
@@ C.1 Simulation Configuration
Through these service requirements, Virne enables the simulation of networks with diverse operational demands, providing users with the flexibility to model complex and dynamic network systems.
@@ C.3.1 RL-based Algorithm Implementations
  • Monte Carlo Tree Search (MCTS) [10] is a planning algorithm that explores the decision space by building a search tree, balancing exploration and exploitation to find action sequences.
  • Policy Gradient (PG or REINFORCE) [64] is a policy-based RL method that learns a parameterized policy to maximize expected returns, i.e., solution quality of NFV-RA.
  • Asynchronous Advantage Actor-Critic (A3C) [31] is an actor-critic-based RL algorithm that uses multiple parallel actors, each interacting with its own copy of the environment. It learns both a policy (actor) and a value function (critic) to estimate the advantage of taking certain actions, leading to more efficient learning than pure PG methods.
  • Proximal Policy Optimization (PPO) [32] is a popular RL method known for its stability, ease of implementation, and strong empirical performance across a wide range of tasks. It achieves policy optimization stability by using a clipped surrogate objective.
  • Deep Q-Network (DQN) and variants [65] are value-based RL methods that aim to learn the state-action value space. However, NFV-RA often involves large, structured action spaces not directly amenable to these methods.
@@
  • Multi-Layer Perceptron (MLP)-based Policy [40, 43, 49, 11]. It concatenates the features of the current to-be-placed virtual node with every physical node’s features. Then, it fed them into multiple fully connected layers and outputs probabilities for selecting each physical node for placement or parameters for action selection.
  • Convolutional Neural Network (CNN)-based Policy [38, 42, 44, 46, 16]. They use CNN to process graph-structured data by treating node features and adjacency information as grid-like inputs. Similar to the MLP, virtual link demands are integrated as node features.
  • Attention-based Policy[26, 47]. It first embeds the features of the current to-be-placed virtual node to form a query. Then, the features of each physical node are similarly embedded to form keys and values. The attention mechanism then computes a weighted sum of physical node values based on their compatibility with the virtual node query.
  • Graph Convolutional Network (GCN)-based policy [27, 45, 14] . It uses GCN to learn node representations by aggregating information from their local neighborhoods. In Virne, GCNs are used to process the PN topology, often with virtual node demands incorporated as additional node features on the PN nodes.
@@
  • GCN & Sequence-to-Sequence (GCN&S2S) Policy [13]. This hybrid architecture combines GCNs for graph representation learning with a Sequence-to-Sequence (S2S) model [66], an RNN-based encoder-decoder with attention, to generate an ordered sequence of actions.
  • GAT&S2S Policy [17]. Similar to GCN&S2S, but replaces the GCN encoder(s) with GAT(s).
  • Graph Attention Network (GAT)-based policy [17]. GATs extend GCNs by incorporating attention mechanisms into the neighborhood aggregation process. Particularly, it also incorporates edge features (e.g., link bandwidth, latency) into the GAT attention mechanism.
  • Dual GCN (DualGCN)-based policy [19]. This architecture uses two separate GCNs, i.e., one to embed the topology and features of the VN, and another for the PN. The embeddings from both graphs are then combined into an MLP to make placement.
  • Dual GAT (DualGAT)-based Policy [19]: Similar to DualGCN, but it utilizes GATs for both VN and PN embedding.
  • Heterogeneous Graph Attention Network (HeteroGAT)-based Policy [39, 50]. Recent studies model NFV-RA instances as a heterogeneous graph, where virtual nodes, physical nodes, virtual links, and physical links are different types of nodes/edges with distinct features. Additionally, HeteroGAT introduces the cross-graph connection links to represent the historical mapping status, i.e, if a virtual node is placed onto a physical node, a heterogeneous link will be created.
@@
C.3.2 Implemented Traditional Algorithms

### v2

@@ A.2 Extensions in Emerging Networks
In modern data center networks, computing resources are often heterogeneous, meaning physical nodes may have varying capacities in terms of CPU, GPU, memory, etc. To account for this, we extend the basic model by incorporating heterogeneous resources, where both virtual and physical nodes are associated with a set of computing resources $\mathcal{C}$. Thus, a physical node $n_{p}$ for placing virtual node $n_{v}$ must have enough resources across all types, i.e., $C(n_{v})\leq C(n_{p}),\forall n_{v}\in\mathcal{N}_{v},C\in\mathcal{C},n_{p}=f_{\mathcal{N}}(n_{v})$.
@@ Appendix B Related Work
As network scenarios become increasingly diverse and complex (e.g., cloud, edge, 5G), NFV has emerged as a critical paradigm for flexible service provisioning. Efficient resource management in NFV spans the entire service lifecycle, encompassing tasks such as efficient allocation (Jin et al., 2020), adaptive scaling (Fei et al., 2018), and dynamic migration (Zhang et al., 2021b). Among these, Resource Allocation in NFV (NFV-RA) forms the foundation. It is typically formulated as an embedding or placement problem, with the central challenge of mapping virtual requirements onto physical infrastructure. While many studies investigate NFV-RA in emerging networks while addressing additional factors (e.g., latency, energy, heterogeneity), their formulations often follow a similar methodological backbone. This shared structure enables a unified perspective on how NFV-RA techniques have evolved over time. In this section, we focus on the methodological aspects and review the development of key approaches in NFV-RA. Then, we describe existing benchmarks to emphasize the gap between Virne. Finally, we compare NFV-RA with other COPs, highlighting the need for specific designs.
@@
(He et al., 2023) Edge / A2C LSTM, Attention Not available
(Zhang et al., 2024a) Edge Energy, Latency PG CNN Not available
(Tan et al., 2024) 5G Slicing Security DQN Heterogeneous GCN Not available
(Liu et al., 2020) 5G Slicing Latency PPO MLP Not available
(Irawan et al., 2023) 5G Slicing Latency PPO GCN Virne
(Tian et al., 2024) Internet of Things Latency PPO GNN Not available
(Fu et al., 2020) Internet of Things Latency DQN CNN Not available
(Guo et al., 2022) Internet of Things - A3C MLP Not available
(Maity et al., 2024) Satellite Latency DQN MLP Not available
(Zhang et al., 2023) Data Plane Latency A3C GCN Not available
(Zhang et al., 2024b) Space-Air-Ground Latency PG CNN Not available
(Haeri and Trajković, 2017) Cloud / MCTS / VNE-Sim
(Dolati et al., 2019) Cloud / PG CNN Not available
(Xiao et al., 2019) Cloud Latency PG MLP Self-implemented
(Solozabal et al., 2020) Cloud Energy PG LSTM; Attention Self-implemented
(Yao et al., 2020) Cloud / PG RNN Not available
(Yan et al., 2020) Cloud / A3C GCN VNE-Sim
(Huang et al., 2021) Cloud / DQN MLP VNE-Sim
(Wang et al., 2021) Cloud / A3C GCN, GRU Virne
(Zhang et al., 2021a) Cloud Security PG CNN Not available
(Ma et al., 2023) Cloud / PG GCN Not available
(Zhang et al., 2022) Cloud / A3C GCN, GRU VNE-Sim
(Geng et al., 2023) Cloud / PPO GNN Virne
(Xiao, 2023) Cloud / PG MLP Not available
(Wang et al., 2023) Cloud / PPO Edge-aware GAT Virne
(Sahraoui et al., 2024) Cloud Energy A2C Attention Virne
(Wang et al., 2024c) Cloud Heterogeneity PPO Cross-GCN Virne
(Wang et al., 2025b) Cloud, Edge Latency PPO Heterogeneous GAT Virne
@@ B.1 Traditional NFV-RA Algorithms
Earlier studies for the NFV-RA problem employed exact solvers, such as integer linear programming (ILP)(Shahriar et al., 2018) and mixed integer programming (MIP)(Chowdhury et al., 2009), to find optimal solutions. However, their high computational complexity makes them impractical for real-world scenarios, especially in larger networks where dynamic service requests require rapid solutions. To address these limitations, heuristic-based methods emerged as alternatives, offering suboptimal yet computationally efficient solutions. Among these, node-ranking strategies stand out as a prominent approach (Gong et al., 2014; Zhang et al., 2018; Fan et al., 2021; 2023). These strategies construct solutions by prioritizing the virtual and physical nodes based on specific metrics to guide the node mapping process, followed by link mapping. For instance, Node Resource Management (NRM) (Zhang et al., 2018) ranks nodes by weighting both node and link resources, while the Node Essentiality Assessment (NEA) (Fan et al., 2023) incorporates topology connectivity into decision-making. Although reducing computational overhead, they rely heavily on manual design with domain-specific knowledge and lack adaptation to diverse scenarios, which can lead to inferior results in complex NFV-RA requirements. Furthermore, meta-heuristics have been adopted to solve NFV-RA by leveraging the evolutionary process (Dehury and Sahoo, 2019; Leivadeas et al., 2013; Fajjari et al., 2011). These algorithms, such as Genetic Algorithms (GA) (Zhang et al., 2019; Dehury and Sahoo, 2019; Wang et al., 2025a) and Particle Swarm Optimization (PSO) (Su et al., 2014; Jiang and Zhang, 2021), explore the solution space by iteratively evolving a population of candidate solutions. However, they are computationally expensive and hyperparameter-sensitive, limiting their practicality.
@@ B.2 RL-based NFV-RA Algorithms
In recent years, machine learning (ML)-based methods have gained prominence in solving NFV-RA problems due to their superior performance and adaptability to dynamic network conditions (Wu et al., 2024a). Several works (Blenk et al., 2016; Thakkar et al., 2020) have explored predictive models trained on high-quality datasets to forecast the future service requests. However, acquiring labeled, high-quality datasets for large-scale and unseen network scenarios remains impractical, limiting their applicability. More dominantly, RL has demonstrated its promise for NFV-RA tasks (Haeri and Trajković, 2017; Xiao et al., 2019; Yao et al., 2020; Wang et al., 2021; Yan et al., 2020; Zhang et al., 2022; Dolati et al., 2019; Wang et al., 2023; Geng et al., 2023; Wang et al., 2024c), mainly due to its label-free nature and adaptation to handle dynamics of network. RL-based methods model NFV-RA as a Markov Decision Process (MDP), allowing an agent to learn optimal policies through iterative interactions with the environment. In general, we unified existing RL-based methods for NFV-RA under a framework with three core components: MDP modeling, policy architecture, and training methods. These methods conceptualize the node mapping allocation process as a sequential decision-making task, where physical nodes are incrementally selected to host VNFs. To execute these decisions, various policy network architectures are designed to represent the network state and generate corresponding actions. Then, they leverage a specific RL method to optimize the policy network with collected experience during interactions. For example, (Xiao et al., 2019) utilized a multilayer perceptron (MLP) and trained it using the policy gradient (PG) algorithm, while the work (Yan et al., 2020) combined MLPs with graph convolutional networks (GCNs) Wang et al. (2024b); Deng et al. (2024) and employed the asynchronous advantage actor-critic (A3C) algorithm. While RL-based methods offer notable advantages in terms of performance and adaptability, there is a lack of in-depth analysis of the impact of specific model design choices and consistent performance comparison. Moreover, systematically identifying the remaining significant challenges in RL-based NFV-RA is critical to open opportunities for future exploration and innovation.
@@ B.3 Existing NFV-RA Benchmarks
For the NFV-RA problem, several benchmarks have been developed for simulation and evaluation. As summarized in Table 1, the most notable existing benchmarks include: VNE-Sim [^1] https://tehreemf.wixsite.com/vne-sim (Haeri and Trajković, 2016), ALEVIN  [^2] https://sourceforge.net/p/alevin/wiki/home/ (Beck et al., 2014), ALib [^3] https://github.com/vnep-approx/alib (Beck et al., 2014), SFCSim [^4] https://github.com/SFCSim/SFCSim (Xu et al., 2022), and Iflye [^5] https://github.com/Echtzeitsysteme/iflye (Tomaszek, 2021). For instance, VNESim(Haeri and Trajković, 2016) supports three heuristic algorithms and focuses exclusively on cloud-based simulations, evaluating solutions based on their effectiveness. Similarly, SFCSim (Xu et al., 2022), incorporates three additional heuristic methods, continuing the trend of focusing on cloud-based scenarios However, most of these benchmarks are limited to cloud-based simulations and lack the flexibility to accommodate more modern network scenarios, such as edge computing or 5G environments. Moreover, these benchmarks generally implement only a few exact and heuristic algorithms and focus mainly on effectiveness evaluation. These limitations highlight the need for more comprehensive benchmarks that can assess algorithms across a broader range of network environments and evaluation perspectives.
@@ B.4 RL for Combinatorial Optimization
RL has gained significant attention for solving Combinatorial Optimization Problems (COPs), where the goal is to learn high-quality solutions (Bengio et al., 2021). These problems span various domains, including the Traveling Salesman Problem (TSP) (Kool et al., 2018), Vehicle Routing Problem (VRP) (Zhou et al., 2023), and bin packing (Zhao et al., 2021). RL-based approaches to solving these problems generally fall into two categories: construction methods that build a solution incrementally from scratch and improvement methods that start with an initial solution and iteratively refine it. Compared with these classic COPs, the NFV-RA problem presents unique challenges, highlighting the need for specialization of RL approach design. First, NFV-RA requires real-time decision-making to meet strict network service requirements. NFV-RA algorithms must execute decisions almost immediately, making construction methods more suitable to incrementally build solutions within time constraints. Second, NFV-RA operates in highly dynamic environments where both service demands and network resources fluctuate in real time. This variability requires RL methods that can adapt to changes in network topologies, resource availability, and demand patterns. Lastly, the state space of NFV-RA is highly complex, involving intricate interactions such as cross-graph mapping and bandwidth-constrained routing. These factors require RL methods to navigate a multidimensional decision space, accounting for diverse resource types and connectivity requirements.
@@ C.1 Simulation Configuration
Through these service requirements, Virne enables the simulation of networks with diverse operational demands. Moreover, Virne supports adaptable constraint modeling to accommodate specific scenarios. For instance, users can enable VNF aggregation strategies (e.g., for SFC) by relaxing node disjointness constraints via the reusable parameter. This provides users with the flexibility to model complex and dynamic network systems beyond standard definitions.
@@ C.2 MDP Variation for NFV-RA
C.2.1 Multi-Task MDP for Meta RL Methods
In practical network systems, VNRs often exhibit significant diversity in their characteristics, such as varying sizes, topologies, and resource demands. A standard single-policy approach struggles to generalize across this wide spectrum. To address this, a multi-task MDP formulation can be employed, framing the embedding of different categories of VNRs as distinct tasks. Specifically, VNRs are grouped into tasks $\mathcal{M}_{i}$ drawn from a distribution $p(\mathcal{M})$. The objective is to train a meta-policy $\pi_{\phi}$ that captures common, transferable strategic knowledge across all tasks. This meta-policy can then be rapidly adapted with minimal data to derive specialized sub-policies $\pi_{\theta_{i}}$ for each task. The meta-optimization objective is to find meta-parameters $\phi$ that maximize the expected performance over the task distribution after adaptation:
$\max_{\phi}\mathbb{E}_{\mathcal{M}_{i}\sim p(\mathcal{M})}\left[J(\theta_{i})\right]=\mathbb{E}_{\mathcal{M}_{i}\sim p(\mathcal{M})}\left[J(f_{\phi}(\mathcal{D}_{i}))\right]$ (9)
where $J(\theta_{i})$ is the expected return for the policy $\pi_{\theta_{i}}$ on task $\mathcal{M}_{i}$, and $\theta_{i}=f_{\phi}(\mathcal{D}_{i})$ represents the parameters of the sub-policy adapted from the meta-policy $\phi$ using a small amount of task-specific experience $\mathcal{D}_{i}$. The meta-policy is then updated by aggregating gradients from the adapted policies, for instance, using a second-order meta-gradient over the task-specific losses.
C.2.2 Constrained MDP for Safe RL Methods
The NFV-RA problem is fundamentally defined by its hard resource constraints, where any violation renders a solution infeasible. Standard MDPs, which typically integrate constraint violations into the reward signal via simple penalties or doing nothing, often fail to guarantee zero-violation solutions. A more rigorous approach is to formulate the problem as a Constrained CMDP. This method explicitly separates the primary reward from costs associated with constraint violations and is formally expressed as a tuple $\langle S,A,P,R,H,C,\gamma\rangle$. Here, in addition to standard MDP components, $H$ is a violation function that measures the degree of constraint violation at each state, and $C$ is a cost function that maps violations to a non-negative cost, i.e., $C(s)=\max(H(s),0)$. The objective is to learn a policy $\pi$ that maximizes the expected cumulative reward $J_{r}(\pi)$ while ensuring the expected cumulative cost $J_{c}(\pi)$ remains below a predefined threshold $d$ (e.g., $d=0$):
$\max_{\pi}$ $J_{r}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{T}\gamma^{t}R(s_{t},a_{t})\right]$ (10)
s.t. $J_{c}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{T}\gamma^{t}C(s_{t})\right]\leq d$
To enforce stricter state-wise safety, this model can be enhanced with reachability analysis. This extension aims to ensure the policy remains within a feasible region where future constraint violations are avoidable, often by optimizing an objective that guarantees the worst-case long-term violation, or feasible value function $V_{h}^{\pi}(s)$, remains non-positive.
@@ C.3.1 RL-based Algorithm Implementations
  • Monte Carlo Tree Search (MCTS) (Haeri and Trajković, 2017) is a planning algorithm that explores the decision space by building a search tree, balancing exploration and exploitation to find action sequences.
  • Policy Gradient (PG or REINFORCE) (Sutton et al., 1999) is a policy-based RL method that learns a parameterized policy to maximize expected returns, i.e., solution quality.
  • Asynchronous Advantage Actor-Critic (A3C) (Mnih et al., 2016) is an actor-critic-based RL algorithm that uses multiple parallel actors, each interacting with its own copy of the environment. It learns both a policy (actor) and a value function (critic) to estimate the advantage of taking certain actions, leading to more efficient learning than pure PG methods.
  • Proximal Policy Optimization (PPO) (Schulman et al., 2017) is a popular RL method known for its stability and strong empirical performance across a wide range of tasks. It achieves policy optimization stability by using a clipped surrogate objective.
  • Deep Q-Network (DQN) and variants (Mnih et al., 2013) are value-based RL methods that aim to learn the state-action value space. However, NFV-RA often involves large, structured action spaces not directly amenable to these methods.
@@
  • Multi-Layer Perceptron (MLP)-based Policy (Liu et al., 2020; Maity et al., 2024; Xiao, 2023; Xiao et al., 2019). It concatenates the features of the current to-be-placed virtual node with every physical node’s features. Then, it fed them into multiple fully connected layers and outputs probabilities for selecting each physical node for placement.
  • Convolutional Neural Network (CNN)-based Policy (Zhang et al., 2024a; Fu et al., 2020; Zhang et al., 2024b; 2021a; Dolati et al., 2019). They use CNN to process graph-structured data by treating node features and adjacency information as grid-like inputs. Similar to the MLP, virtual link demands are integrated as node features.
  • Attention-based Policy(He et al., 2023; Solozabal et al., 2020). It first embeds the features of the current to-be-placed virtual node to form a query. Then, the features of each physical node are similarly embedded to form keys and values. The attention mechanism then computes a weighted sum of physical node values based on their compatibility with the virtual node query.
  • Graph Convolutional Network (GCN)-based policy (Zhang et al., 2023; Ma et al., 2023; Yan et al., 2020) . It uses GCN to learn node representations by aggregating information from their local neighborhoods. In Virne, GCNs are used to process the PN topology, often with virtual node demands incorporated as additional node features on the PN nodes.
@@
  • GCN & Sequence-to-Sequence (GCN&S2S) Policy (Wang et al., 2021). This hybrid architecture combines GCNs for graph representation learning with a Sequence-to-Sequence (S2S) model (Sutskever et al., 2014), an RNN-based encoder-decoder with attention, to generate an ordered sequence of actions.
  • GAT&S2S Policy (Wang et al., 2023). Similar to GCN&S2S, but replaces the GCN encoder(s) with GAT(s).
  • Graph Attention Network (GAT)-based policy (Wang et al., 2023). GATs extend GCNs by incorporating attention mechanisms into the neighborhood aggregation process. Particularly, it also incorporates edge features (e.g., link bandwidth, latency) into the attention calculation.
  • Dual GCN (DualGCN)-based policy (Wang et al., 2024c). This architecture uses two separate GCNs, i.e., one to embed the topology and features of the VN, and another for the PN. The embeddings from both graphs are then combined into an MLP to make placement.
  • Dual GAT (DualGAT)-based Policy (Wang et al., 2024c): Similar to DualGCN, but it utilizes GATs for both VN and PN embedding.
  • Heterogeneous Graph Attention Network (HeteroGAT)-based Policy (Tan et al., 2024; Wang et al., 2025b). Recent studies model NFV-RA instances as a heterogeneous graph, where virtual nodes, physical nodes, virtual links, and physical links are different types of nodes/edges with distinct features. Additionally, HeteroGAT introduces the cross-graph connection links to represent the historical mapping status, i.e, if a virtual node is placed onto a physical node, a heterogeneous link will be created.
@@
C.3.2 Traditional Algorithm Implementation